How can I solve this equation -(-53-65)-(55+3)=55
Answer by CharlesG2(834)

How can I solve this equation -(-53-65)-(55+3)=55
-(-53-65)-(55+3)=55
53+65-55-3=55 (carried through the -1)
118-55-3=55
63-3=55
60=55, hmmm 60 does not equal 55, there is something wrong here